Stats help!!

A researcher wants to determine if the distribution of asthma mortality by race has changed over the past years. He gathered data from Center of Disease Control and Prevention, and the following data was obtained from National Asthma Mortality Rate of 2015:

Race White Non-Hispanic Black Non-Hispanic Other Non-Hispanic Hispanic Total
Observed Counts (O) 2259 1148 235 358 4000
Percentages 60% 25% 5% 10% 100%
Expected Counts (E)

1) If the null hypothesis is true, what are the expected counts?

A) E1=60, E2=25, E3=5, E4=10

B) E1=2400, E2=1000, E3=200, E4=400

C) E1=1355.4, E2=287, E3=11.75, E4=35.8

D) E1= E2= E3 = E4= 1000

E) E1=2259, E2=1148, E3=235, E4=358

2) The appropriate null hypothesis for the test is:

A) At least one proportion is different from what is expected

B) P1=2259, P2=1148, P3=235, P4=358

C) P1=2259/4000, P2=1148/4000, P3=235/4000, P4=358/4000

D) P1=0.6, P2=0.25, P3=0.05, P4=0.1

E) P1=60, P2=25, P3=5, P4=10
Answer #1

1) Total is 4000

Expected value for White Non-Hispanic = 0.6*4000 = 2400

Expected value for Black Non-Hispanic = 0.25*4000 = 1000

Expected value of other non-hispanic = 0.05*4000 = 200

Expected value of hispanic = 0.1*4000 = 400

Hence, Option B) E1 = 2400, E2 = 1000, E3 = 200, E4 = 400 is the correction option for this.

2) Option D) P1 = 0.6, P2 = 0.25, P3 = 0.05, P4 = 0.1 is the correct answer for this.